OpenStudy (anonymous):

Simplify the radical expression. All variables represent positive real numbers. square root of 700x squared y

OpenStudy (stamp):

\[\sqrt700x^2y\]

OpenStudy (stamp):

What jumps out at you right away?

OpenStudy (anonymous):

the \[x^{2}\]

OpenStudy (stamp):

\[\sqrt (x^2)\sqrt700\sqrt y\]\[x\sqrt700y\]Can we go further?

OpenStudy (anonymous):

10xy\[\sqrt{7}\]?

OpenStudy (stamp):

We cannot pull out the y, but good job in identifying the 100 = 10(10)\[x\sqrt (7^110^2y^1)\]\[10x\sqrt7y\]

OpenStudy (anonymous):

thanks :)